A solution contains 5.0 g ethylene glycol, C2H4(OH)2, and 95 g water, H2O, by mass. The density of the solution is 1 g/mL. Find the a) freezing point of the solution b) osmotic pressure at 35°C.


1
Expert's answer
2021-01-15T07:01:36-0500

a) The freezing point depression:

"∆T = k_f \times c_m"

kf = the molal freezing point depression constant

cm = the molality of the solute

M(C2H4(OH)2) = 62.06 g/mol

n(C2H4(OH)2) "= \frac{5.0}{62.06} = 0.0805 \;mol"

c_m = 0.805 M

k_f = 3.11 (from table for ethylene glycol)

∆T = "3.11 \times 0.805 = 2.50" ºC

The freezing point of the solution = 0 ºC – 2.50 ºC = -2.50 ºC

b) T = 35 + 273 = 308 K

Osmotic pressure:

π = MRT

M = molality of the solution

R = gas constant

T = temperature

"π= 0.805 \times 0.082 \times 308 = 20.33 \;atm"